OpenStudy (anonymous):

if sin ( x + 20°) = cos x, then the value of x is

OpenStudy (anonymous):

sin = cos at 45o angle x + 20 = 45, x = 25.

OpenStudy (campbell_st):

well this is how the question is done cos(x) = sin(90 - x) this it the complementary angles and trig ratios property here is an example  but anyway you can now say \[\sin(x + 20) = \sin(90 - x)\] since the ratios are the same you can equate the angles x + 20 = 70 - x just solve for x... hope this helps... p.s. the answer isn't 25 degrees